OpenStudy (kaylaprincess):

I dont understand :\ What is NOT a restriction of the product of quantity x squared plus 3 x plus 2 divided by quantity x squared minus 2 x minus 3 multiplied by quantity x squared plus 4 x pl ? (2 points)

OpenStudy (kaylaprincess):

\[ \frac{ 3x + 12x + 9 }{ x^2 - 4 } \div \frac{ 4x + 4 }{ x + 2 }\]

OpenStudy (anonymous):

Because this is division, you have to flip the 2nd fraction and change it to multiplication. \[\frac{ 3x + 12x + 9 }{ x^2 - 4 } \times \frac{ x+2 }{ 4x+4 }\] You have to solve for the restrictions to find numbers that are NOT restrictions. The restrictions are any numbers that make the denominator 0. So you have to solve these two equations to get the restrictions \(x^2-4=0\) and \(4x+4=0\)

OpenStudy (kaylaprincess):

would it be -4? as the restriction?

OpenStudy (anonymous):

no, you can always test your options in the denominators. If you get 0, it's a restriction

OpenStudy (kaylaprincess):

i still feel confused.

OpenStudy (kaylaprincess):

does the number have to be same to be the restriction to both?

OpenStudy (anonymous):

not necessarily. each fraction will have it's own restriction. The restrictions for the whole problem will be all of them combined

OpenStudy (anonymous):